### v1

# Finitely many worlds with finite information flow and Bell theorem

Abstract
The Copenhagen interpretation of quantum theory does not provide a measurement-independent description of physical processes. Two competing approaches to this issue are single-world ontological theories and Everett many-worlds theory. The former addresses the measurement problem by introducing auxiliary random variables that specify, in particular, the actual values of macroscopic observables. Such models may be $\psi$-epistemic, meaning the quantum state is not part of the ontology. However, all the ontological models break the principle of locality, as demonstrated by Bell’s theorem. The latter approach, which is deterministic and $\psi$-ontic, offers the advantage of evading Bell’s theorem by delaying the buildup of the correlations until the parties compare their outcomes at a meeting point. Because of determinism and the ontic nature of the quantum state, the number of parallel worlds and the information flow towards the meeting point turn out to be infinite. By integrating the strengths of the two approaches, we introduce a simple $\psi$-epistemic, many-worlds, local model of projective measurements on two spatially separate maximally entangled qubits. Because of its randomness, the model requires only two branches and a finite information flow – just one bit per measurement is communicated to the meeting point. We explore how this hybrid approach, employing both randomness and branching, addresses key challenges of each individual framework.
.0.1 Introduction
The double-slit and Stern-Gerlach experiments are classic textbook examples that illustrate the core puzzle of quantum theory, that is, the superposition and inteference of distinguishable alternative states. According to the Copenhagen interpretation, as long as a system is in a superposition of two or more states, none of them is actualized. Schrödinger famously illustrated the extreme implications of this view with his thought experiment, in which a cat is placed in a superposition of being both dead and alive trimmer . This scenario highlights a fundamental tension between the interpretation and our intuitive expectation that the cat must experience one definite state.
In modern terms, the apparent paradox in Schrödinger’s thought experiment finds a pragmatic resolution in decoherence theory. Since no experiment can detect interference between the ”dead” and ”alive” states, the superposition becomes de facto equivalent to a statistical mixture of the two. This allows for a consistent reduction to one of the two states without leading to contradictions. Thus, it is safe to claim that one of the alternatives is actualized with no future disproval of the claim.
While this approach leads to an interpretation of quantum theory which is practically consistent, it also raises a conceptual problem: how can one state in a superposition be actualized merely on the promise that no future observation will reveal the superposition? It would seem that the system has a kind of foresight of the future. For example, the system could decohere by emitting a photon going into deep space. Since it is plausible that the photon is lost forever, the system could safely ‘decide’ to reduce its state. However how does it know that the photon will not encounter a mirror on one of Jupiter’s moons and be reflected back for reabsorption?
Since the inception of quantum theory, interpretative challenges have raised questions about its completeness epr . To reconcile state superpositions with the definiteness of the our macroscopic experience, a minimal requirement is the existence of auxiliary information that determines the actual macroscopic state that is experienced. This idea is pursued in ‘hidden-variable’ theories and is consistently realized in the de Broglie-Bohm (dBB) theory bohm . In the dBB theory, both the wave-function and the position of the particles specify the actual state of a system. Employing a term coined in the context of quantum foundations, the state represented by the wave-function and the auxiliary variables is referred as an ontic state. More broadly, instead of variables, the precise ontic state itself could be hidden. Therefore, more generally, we refer to these hypothetical completions as ontological theories.
It is important to emphasize that ontological theories are inherently statistical, reflecting the randomness of quantum phenomena. In Bohmian theory, for instance, particle positions are randomly distributed according to Born’s rule. As we will see, however, the randomness of a theory of quantum phenomena is not self-evident.
Specific ontological theories, known as $\psi$-epistemic theories, has gained significant interest over the past decade pbr ; hardy ; leifer0 ; barrett ; leifer ; ring . In this subclass, the quantum state is interpreted as a state of knowledge and is not part of the ontology; instead, its information is encoded in the statistics of the ontic state. This statistical encoding provides a straightforward explanation of why quantum states cannot be cloned or why two non-orthogonal states cannot be discriminated with certainty. Furthermore, there is a link between $\psi$-epistemic theories and classical simulations of quantum communication which employ finite classical communication montina_comm_epist0 ; montina_comm_epist .
However, this route toward a realistic completion of quantum theory leads to a series of conceptual problems, such as contextuality kochen and, more seriously, nonlocality bell , on which we focus in this paper. As shown by Bell, the correlations between two entangled systems are incompatible with any (single-world) local ontological theory; a measurement performed on one system would instantaneously affect the ontic state of the other, in apparent violation of the principle of locality.
Another approach to addressing the interpretative problems of quantum theory is the Everett many-worlds theory everett . Since a cat must be either dead or alive after all, Everett proposed that both states are equally realized, but in separate worlds. This accounts for extremely unlikely yet theoretically possible future interferences between the alternatives. Every time a system enters a superposition of states, it results in a branching into parallel worlds, leading to an exponential blowup of distinct realities over time.
What this theory preserves are the relations between events within each world. However, a fundamental issue arises: where do probabilities come from? Specifically, the theory does not inherently explain why certain events are more probable than others. For example, consider a system evolving into the superposition
$\psi_{+}|+\rangle+\psi_{-}|-\rangle$ (1)
[…]
A possible resolution of this issue is to assume that there is a multitude of branches such that the ratio between the branches in |+ket|$\rangle$| + ⟩ and those in |ket|$\rangle$| - ⟩ is about $|\psi_{+}/\psi_{-}|^{2}$ saunders . Since this ratio is a real number, the Born rule is recovered exactly in the limit of infinite branches. Thus, if a system splits into a finite number of macroscopically distinct states, each one actually corresponds to an infinite multitude of parallel worlds. Hereafter, in the context of many-worlds theory, we employ this derivation of Born rule.
While (single-world) ontological theories are probabilistic, the many-worlds theory is $\psi$-ontic and deterministic. Randomness arises from the subjective perspective of an observer following a single branch of the tree. In this paper, we explore how integrating the randomness of ontological theories with many-worlds branching can yield a more economical framework while preserving an advantage of many-worlds theory – its claimed ability to provide a local, realistic picture of the world(s) deutsch . The argument supporting this ability is clearly illustrated by a simple local realistic model of a ‘nonlocal’ Popescu-Rohrlich (PR) box brassard . The core idea of the argument is to delay the buildup of correlations until the parties compare their outcomes at a meeting point. However, many-worlds theory requires an infinite number of branching and an infinite information flow to the meeting point deutsch .
In this paper, we show that just one classical bit of information about the chosen measurement and two branches are sufficient to simulate local projective measurements on maximally entangled qubits. This drastic reduction of information is a consequence of the probabilistic and $\psi$-epistemic nature of the proposed model. While in the many-worlds theory quantum states are part of the ontology, which must therefore contain infinite classical information, in our model the quantum state is rather encoded in the statistics of the ontic states.
Our model directly builds on the Toner-Bacon model, which simulates maximally entangled states using one bit of communication and shared randomness toner . By selecting a branch at random, the model reproduces the same statistics as the single-world model in Ref. gisin , in which outcomes are generated through shared randomness and the single use of a nonlocal PR box.
.0.2 Nonlocal correlations in many-worlds theory
In Ref. deutsch , Deutsch and Hayden proposed a many-worlds, local description of nonlocal quantum correlations using the Heisenberg picture. Here, we present a highly simplified caricature of their argument. While this oversimplification may make the argument seem somewhat trivial, it nevertheless captures its essence.
Suppose that two qubits are prepared in the singlet state
$|\Psi\rangle=\frac{1}{\sqrt{2}}\left(|-1,-1\rangle-|1,1\rangle\right).$
The qubits are sent to two spatially separated parties, Alice and Bob. Each party independently chooses to perform a projective measurement, represented by the Bloch vectors $\vec{a}$ and $\vec{b}$, respectively. The outcomes, ±$\pm 1$, of each measurement occur with equal probability. In the MW framework, this means that qubits locally branch into a multitude of realities in half of which a party observes one of the two outcomes. On each side, this process occurs without the need to know what is happening on the other side. After the measurements, the parties decide to compare their results. To do that, they have to meet at the same place or send some communication through a classical channel. After multiple running of the experiment over different entangled states, they find out that the outcomes, $s_{A}\in\{-1,1\}$ and $a_{B}\in\{-1,1\}$, are distributed according to the probability distribution
$P(s_{A},s_{B}|\vec{a},\vec{b})=\frac{1}{2}\left(1-s_{A}s_{B}\vec{a}\cdot\vec{b}\right),$ (2)
which violate the Bell inequalities. While this violation leads to a break of local realism in a single-world scenario, the MW framework evades this conclusion as follows. When Alice meets with Bob, the multitudes of realities of each party are paired in such a way that the fraction of realities experiencing the outcomes $s_{A}$ and $s_{B}$ is exactly equal to $P(s_{A},s_{B}|\vec{a},\vec{b})$. This argument is illustrated in Ref. brassard for nonlocal PR boxes and further discussed in the general quantum case in Ref. charles . The existence of many local realities allows us to delay the buildup of the correlations until the parties meet or communicate each other.
It is fundamental to remark that, in the MW theory, each party has to carry the information about the performed measurement at the meeting point in order to get the right pairing. This information is clearly infinite, which is a consequence of the $\psi$-ontic nature of the theory.
.0.3 Finitely many worlds with finite information flow
Now, let us show that two branches and a finite information flow are enough for simulating the outcomes of all the projective measurements $\vec{a}$ and $\vec{b}$ on a maximally entangled state. The proof is a straightful application of MW argument to Toner-Bacon model toner . The model provides a classical simulation of the outcomes by using one bit of communication between the parties and some shared randomness.
The classical protocol is as follows. Alice and Bob share two random and independent unit vectors, $\vec{x}_{0}$ and $\vec{x}_{1}$, uniformly distributed over a sphere. First, Alice chooses to simulate the measurement $\vec{a}$ and generates an outcome $s_{A}$ such that
$s_{A}=\text{sign}(\vec{a}\cdot\vec{x}_{0}).$ (3)
Then, she generates a number
$n_{A}=\text{sign}(\vec{a}\cdot\vec{x}_{0})\text{sign}(\vec{a}\cdot\vec{x}_{1}),$ (4)
which is sent to Bob. Finally, Bob chooses to simulate the measurements $\vec{b}$ and generates the outcome
$s_{B}=-\text{sign}\left[\vec{b}\cdot(\vec{x}_{0}+n_{A}\vec{x}_{1})\right].$ (5)
Averaging on $\vec{x}_{0}$ and $\vec{x}_{1}$, it turns out that the outcomes are generated according to Eq. (2). Let us prove that. It is sufficient to show that
$\langle s_{A}s_{B}\rangle=-\vec{a}\cdot\vec{b}.$ (6)
We have
$\begin{array}[]{l}\langle s_{A}s_{B}\rangle=\frac{1}{(4\pi)^{2}}\sum_{n_{A}=\pm 1}\int d^{2}x_{0}d^{2}x_{1}\text{sign}(\vec{a}\cdot\vec{x}_{0})\vspace{1mm}\\ \text{sign}\left[-\vec{b}\cdot(\vec{x}_{0}+n_{A}\vec{x}_{1})\right]\frac{1+n_{A}\text{sign}(\vec{a}\cdot\vec{x}_{0})(\vec{a}\cdot\vec{x}_{1})}{2}.\end{array}$ (7)
First, we note that the two terms of the sum over $n_{A}$ are identical (replace $\vec{x}_{1}$ with $-\vec{x}_{1}$ in the second term). Using the distributive property and a suitable swapping $\vec{x}_{0}\leftrightarrow\vec{x}_{1}$ , we have
$\langle s_{A}s_{B}\rangle=\frac{1}{8\pi^{2}}\int d^{2}x_{0}d^{2}x_{1}\text{sign}(\vec{a}\cdot\vec{x}_{0})\text{sign}\left[-\vec{b}\cdot(\vec{x}_{0}+\vec{x}_{1})\right].$ (8)
Integrating over $\vec{x}_{1}$ in spherical coordinates with $\vec{b}$ as the pole, we have
$\langle s_{A}s_{B}\rangle=\frac{1}{2\pi}\int d^{2}x_{0}(-\vec{b}\cdot\vec{x}_{0})\text{sign}(\vec{a}\cdot\vec{x}_{0}),$ (9)
which can be recast in the form
$\langle s_{A}s_{B}\rangle=\frac{1}{\pi}\int d^{2}x_{0}(-\vec{b}\cdot\vec{x}_{0})\theta(\vec{b}\cdot\vec{x}_{0})\theta(\vec{a}\cdot\vec{x}_{0})-(\vec{a}\rightarrow-\vec{a})$ (10)
where $\theta(x)$ is the Heaviside function. The first term appears in the Kochen-Specker model kochen and is $-(1+\vec{a}\cdot\vec{b})/2$. Thus, Eq. (6) is proved.
Now, let us reinterpret this model within a many-worlds framework, where each party branches into two distinct alternatives, labeled $A_{\pm 1}$ and $B_{\pm 1}$ for Alice and Bob, respectively. First, we note that the Toner-Bacon model generates the same correlation if $s_{A}$ and $s_{B}$ are both chosen with opposite sign. Let us assume that alternatives $A_{1}$ and $A_{-1}$ get the outcome $s_{A}$ in Eq. (3) and its opposite, respectively. They both generate $n_{A}$ according to Eq. (4). All these processes involve only local information that is available to Alice. A complication arises in deciding which outcomes are associated with Bob’s alternatives. Associating $s_{B}$ in Eq. (5) to alternative $B_{1}$ would require to know the value $n_{A}$, an information that is not available to Bob. If we set $n_{A}$ equal to $1$ and, finally, we pair alternative $A_{w}$ with $B_{w}$ at the meeting point, we end up making a mistake in the case $n_{A}$ were $-1$ and $\text{sign}\left[\vec{b}\cdot\vec{x}_{+}\right]\neq\text{sign}\left[\vec{b}\cdot\vec{x}_{-}\right]$, where $\vec{x}_{\pm}\equiv\vec{x}_{0}\pm\vec{x}_{1}$. Crucially, the mistake can be corrected by knowing just one classical bit of information per measurement, namely $n_{A}$ for Alice’s measurement and
$n_{B}\equiv\text{sign}(\vec{b}\cdot\vec{x}_{+})\text{sign}(\vec{b}\cdot\vec{x}_{-})$ (11)
for Bob’s measurement. Thus, let us assume that alternative $B_{1}$ and $B_{-1}$ generate the outcomes
$s_{B}=-\text{sign}\left[\vec{b}\cdot(\vec{x}_{0}+\vec{x}_{1})\right]$ (12)
and $-s_{B}$, respectively. Furthermore, both Bob’s alternatives generate $n_{B}$ according to Eq. (11), which will be used to correct the pairing. Finally, Bob and Alice reach the meeting point to compare their results. Bob and Alice carry information about their outcome and the numbers $n_{A}$ and $n_{B}$, respectively. Alternatives $A_{\pm 1}$ are paired with $B_{\pm 1}$, unless $n_{A}=n_{B}=-1$, in which case the pairing is swapped. This leads to the following pairing rule
$\begin{array}[]{c}(n_{A},n_{B})\neq(-1,-1)\Rightarrow A_{\pm 1}\leftrightarrow B_{\pm 1}\\ (n_{A},n_{B})=(-1,-1)\Rightarrow A_{\pm 1}\leftrightarrow B_{\mp 1}.\end{array}$ (13)
This model is mathematically equivalent to the model in Ref. gisin , but they differ in the interpretation. The latter is a single-world model in which the parties share a nonlocal PR box as a resource for generating the outcome. In our model, we have a ”local” PR box which is used for pairing locally the alternatives at the meeting point. The statistics of the single-world model of Ref. gisin is recovered by randomly taking one of the two paired alternatives in the ”many-worlds” model.
It is interesting to note that a local single-world model could be constructed in a slightly different scenario. Instead of generating the correct outcomes at the moment measurements are performed, the task shifts to generating them at the meeting point. Each party only needs to communicate two bits. However, including the whole process in this local description would lead to the paradoxical conclusion that a party’s state – and even their memory of past observations – can suddenly change. In a many-worlds framework, this macroscopic ‘rewind’ and memory reconstruction is avoided by allowing two distinct realities to evolve independently.
.0.4 Discussion
The nonlocality arising in the framework of ontological theories can lead the proponents of many-worlds theory to claim the failure of a probabilistic framework. The main purpose of this paper is to show with a proof of principle that only the branching of the worlds is required for a local account of the correlations of two maximally entangled qubits. There is no known proof that both determinism and the ontology of the quantum states are necessary conditions for a local theory. Furthermore, the drop of these two assumptions can lead to a more ‘economical’ theory, in which the information flow and the number of branches is made as small as possible.
[…]
The second difference is more subtle. In a $\psi$-epistemic theory, the ontic state does not contain the full information about the quantum state. Rather, this information is encoded in the statistics of the ontic state. Let us consider the quantum state emerging from Alice’s measurement and how its information is encoded in many-worlds theory and our model. In the former, the ongoing state, which has to be carried to the meeting point, is the vector $\vec{a}$ and the outcome associated with each branch. Their knowledge allows us to infer the outgoing quantum state after the measurement in each branch, meaning that the quantum state is part of the ontology. In the latter, this inference is not possible in a single run of the experiment. After the measurement, Alice generates a number $n_{A}$ and associates an outcome to each branch. This information and the knowledge of the shared random variables $\vec{x}_{0}$ and $\vec{x}_{1}$ enable us only to infer that the quantum state $\vec{a}$ in branch $A_{1}$ after the measurement is any unit vector such that Eqs. (3,4) are fulfilled. Rather, the full information about quantum state is statistically encoded in the ontic variables. This is necessary for making the information flow from the parties to the meeting point finite.
.0.5 Conclusions
In conclusion, a hybrid approach that combines the randomness of ontological theories with the branching structure of many-worlds theory can address the challenges of each individual framework. On one hand, branching allows us to evade Bell’s theorem, which applies to single-world ontological theories. More broadly, it may also circumvent the contextuality requirement. For instance, a manifestation of preparation contextuality is the proof that ontological theories inherently break time symmetry leifer1 . However, with appropriate adaptations, our model provides a fully time-symmetric description of a scenario in which a qubit, in a maximally mixed state, is measured at two different times. The time asymmetry proved in Ref. leifer1 is transferred to the measurement devices and the subsequent comparison of results, which are inherently time-asymmetric processes. Similarly, the branching circumvents the recently proved theorem that measurements erase information monti_erase , which is another manifestation of contextuality. On the other hand, the introduction of randomness in the many-worlds framework can make it more natural and economical to account for the observed randomness by allowing us to reduce the number of branches and the information flow as much as possible – it may be possible that there is no infinite branching and not all the alternatives in a superposition are necessarily realized. If there were some branch with a very faint amplitude in which Scipio was defeated in the Third Punic War, this would not necessarily imply that there is a world in which Carthage rules over Rome!
Here, we have presented a proof of principle of this hybridation by showing that local measurements on two maximally entangled qubits can be locally simulated with only two branches, finite communication and shared randomness. It is plausible that a generalization to many qubits and any quantum state will require a blow-up of the shared variables. This increase of local resources occurs also in the local version of the many-worlds theory deutsch , in which the local observables are represented by matrices acting on the whole Hilbert space of all involved particles. The main difference, however, is that the shared random variables in our approach would not depend on the state of the system – they would be generated before the ”game” starts and would provide a stochastic background over which the system evolves. A similar idea has been suggested in Ref. montina_non_markov within a single-world scenario. However, whether this hybrid approach can be extended to a general theory remains an open and, in our opinion, fascinating question.
.0.6 Acknowledgments
We wish to thank Charles Bédard for useful comments and suggestions. The authors acknowledge support from the Swiss National Science Foundation (SNSF), project No. 200020- 214808.