### v3

# Local theories with parallel realities and the epistemic view of the quantum state

Abstract
Hidden-variable theories, such as the de Broglie-Bohm theory, address the measurement problem by introducing auxiliary random variables that specify, in particular, the actual values of macroscopic observables. Such models may be $\psi$-epistemic, meaning the quantum state merely represents a state of knowledge, as opposed to $\psi$-ontic models, in which the quantum state is a physical variable of the theory. A serious issue of this route towards a realistic completion of quantum theory is raised by Bell’s proof that the resulting theories are nonlocal. A possible resolution is to reject the assumption that measurements have single actual outcomes. Indeed, relaxing this premise, Deutsch and Hayden showed that Bell’s theorem can be evaded by delaying the buildup of the correlations until the parties compare their outcomes at a meeting point. However, the Deutsch-Hayden theory, which is deterministic and $\psi$-ontic, leads to an infinite information flow towards the meeting point. Furthermore, alternative branches are weighted by amplitudes, leading to interpretative issues. In this paper, we introduce a general framework that combines the randomness of single-world theories with the coexistence of diverse instances, as found in many-worlds theory. This framework incorporates the existing theories as limiting cases. We explore how this hybrid approach addresses key challenges of single-world and Deutsch-Hayden theories. On one hand, the multiplicity of coexisting instances allows us to circumvent nonlocality and, possibly, contextuality. On the other hand, randomness makes it possible to derive quantum probabilities from unweighted counts of instances and ensemble averages. Furthermore, it can lead to a reduction of the information flow by stripping the quantum state of its ‘ontic’ rank. We illustrate this framework with a simple $\psi$-epistemic local model of projective measurements on two spatially separate maximally entangled qubits. Because of its randomness, the model requires two unweighted instances and a finite information flow – just one bit per measurement is communicated to the meeting point. Setting aside its foundational motivations, this framework has also relevance in quantum communication complexity and leads to novel technical questions, potentially providing new insights into some peculiarities of entanglement.
I Introduction
The double-slit and Stern-Gerlach experiments are classic textbook examples that illustrate the core puzzle of quantum theory, that is, the superposition and interference of distinguishable alternative states. According to the Copenhagen interpretation, as long as a system is in a superposition of two or more states, none of them is actualized. Schrödinger famously illustrated the extreme implications of this view with his thought experiment, in which a cat is placed in a superposition of being both dead and alive trimmer . This scenario highlights a fundamental tension between the interpretation and our intuitive expectation that the cat must experience one definite state.
In modern terms, the apparent paradox in Schrödinger’s thought experiment finds a pragmatic resolution in decoherence theory. Since no experiment can detect interference between the ”dead” and ”alive” states, the superposition becomes de facto equivalent to a statistical mixture of the two. This allows for a consistent reduction to one of the two states without leading to contradictions. Thus, it is safe to claim that one of the alternatives is realized with no future disproof of the claim.
While this approach leads to an interpretation of quantum theory which is practically consistent, it also raises a conceptual problem: how can one state in a superposition be actualized merely on the promise that no future observation will reveal the superposition? It would seem that the system has a kind of foresight of the future. For example, the system could decohere by emitting a photon going into deep space. Since it is plausible that the photon is lost forever, the system could safely ‘decide’ to reduce its state. However how does it know that the photon will not encounter a mirror on one of Jupiter’s moons and be reflected back for re-absorption?
Since the inception of quantum theory, interpretative challenges have raised questions about its completeness epr . To reconcile state superpositions with the definiteness of our unique macroscopic experience, a minimal requirement is the existence of auxiliary information that determines the actual macroscopic state that is experienced. This idea is pursued in “hidden-variable” theories and is consistently realized in the de Broglie-Bohm (dBB) theory bohm . In the dBB theory, both the wave function and the position of the particles specify the physical state of a system. Employing a term coined in the context of quantum foundations, the state represented by the wave function and the auxiliary variables is referred to as an ontic state, since it represents the ontology of the theory. Generally, we refer to these hypothetical completions as ontological theories. In particular, theories in which the wave function is considered a physically real object are called $\psi$-ontic. See Ref. harrigan for a comprehensive discussion on this classification. It is important to emphasize that ontological theories are inherently probabilistic, reflecting the randomness and unpredictability of quantum phenomena. In Bohmian theory, for instance, particle positions are randomly distributed according to Born’s rule.
However, this route towards a realistic completion of quantum theory leads to a series of conceptual problems, such as contextuality kochen and, more seriously, nonlocality bell , on which we focus in this paper. As shown by Bell, the correlations between two entangled systems are incompatible with any single-world local ontological theory; a measurement performed on one system would instantaneously affect the ontic state of the other, in apparent violation of the principle of locality.
The assumption of a single actual outcome of a measurement is an essential hypothesis of Bell’s theorem. Indeed, Deutsch and Hayden (DH) developed a local many-worlds theory deutsch based on Everett’s interpretation everett . In the latter, the wave function is identified as the complete ontology of quantum theory, so that its branching into alternative macroscopically distinct states is a physically real process. Thus, while single-world ontological theories are probabilistic, the many-worlds theory is $\psi$-ontic and deterministic. Randomness arises from the subjective perspective of an observer following a single branch of the tree. By employing the Heisenberg picture, Deutsch and Hayden introduced additional ontological objects – called descriptors – in their local theory. Their argument for locality is clearly illustrated through a simple local realistic model of a ‘nonlocal’ Popescu-Rohrlich (PR) box brassard . The core idea of the argument is to delay the buildup of correlations until the parties compare their outcomes at a meeting point. This delay is possible only if multiple realities coexist locally.
However, since DH theory is $\psi$-ontic, there is an infinite flow of information from the parties to the meeting point. Furthermore, due to determinism, the theory faces the same issue as Everett’s: where do quantum probabilities come from? Although there is extensive literature on the derivation of Born’s rule, no broad consensus exists on how quantum probabilities emerge in Everett’s interpretation. Without entering into this debate, we simply assume – within the context of the many-worlds theory – that branches are weighted by an amplitude, and our expectation of being in a particular branch is given by Born’s rule, which we adopt as a postulate. Occasionally, we will also consider the branch counting proposed by Saunders saunders , since this has some similarity with the counting that we will employ in the framework discussed in this work. Motivated by both foundational considerations and communication-complexity questions, our aim is to investigate alternative local theories that employ fewer resources and unweighted counting rules. To this end, we introduce a general framework that encompasses both single-world and many-worlds theories as limiting cases.
In the search for an alternative local theory with multiple instances, it is fundamental to emphasize that the branching of the wave function constitutes a physical process in both the dBB and Everett theories. In the dBB theory, however, the positions of the particles determine which branch is actually experienced. For this reason, some proponents of the many-worlds interpretation argue that dBB is effectively a many-worlds theory augmented with additional (and, in their view, unnecessary) variables. See, for example, the discussion in Ref. brown and references therein. Indeed, it may seem peculiar that alternative branches of the wave function – each containing detailed biophysical processes, such as brain activity – exist physically, yet are empty of observers. These brain activities in “empty” branches would be akin to “philosophical zombies” introduced in philosophy of mind and in debates about consciousness.
With this important premise on the dBB theory, one might wonder whether any completion of quantum theory is essentially just a many-worlds theory augmented with extra-variables. The crucial point of the question lies in the interpretation of the wave function. If the wave function is regarded as part of the ontology, then it makes sense to claim that the various branches of the wave function correspond to physically existing realities, even if observers, attached to the extra-variables, may follow only one of them. Conversely, each branch of the wave function does not necessarily have physical reality in specific ontological theories, known as $\psi$-epistemic theories, which have gained significant interest over the past decade pbr ; hardy ; leifer0 ; barrett ; leifer ; ring . In this subclass, the quantum state is interpreted as a state of knowledge and is not part of the ontology; instead, its information is encoded in the statistics of the ontic state. Because of this encoding, $\psi$-epistemic theories are essentially equivalent to classical simulations of quantum communication which employ finite classical communication montina_comm_epist0 ; montina_comm_epist . In a $\psi$-epistemic ontological theory, the coexistence of multiple realities would not be justified by the branching of the wave function, but rather by Bell’s theorem. Furthermore, in such theories, the branching of the wave function would not necessarily imply a branching into distinct realities. Since $\psi$-epistemic ontological theories are the only alternative to $\psi$-ontic many-worlds theories with possible extra-variables, we focus on them. An exception worth noting is the many interacting worlds (MIW) approach of Ref. hall , which is $\psi$-ontic in some sense, as discussed later. However, it is not equivalent to many-worlds theory, since the equivalence to quantum mechanics holds only in the limit of infinite worlds (an interesting question regarding this approach will be discussed in the conclusions).
In this paper, we explore whether integrating the randomness inherent in single-world ontological theories with the coexistence of multiple instances – as in the many-worlds theory – can yield a more general framework, while preserving the locality of the Deutsch–Hayden (DH) theory. In particular, the introduction of randomness is a prerequisite for a finite information flow, which is infinite in the DH theory. Moreover, relaxing the determinism of the many-worlds theory makes it possible to derive quantum probabilities through a simple unweighted count of instances and an ensemble average. We illustrate this hybrid framework that incorporates both randomness and multiplicity of instances by showing that just one classical bit of information about the chosen measurement and two unweighted instances are sufficient to simulate local projective measurements on maximally entangled qubits. This model raises the question whether a local $\psi$-epistemic theory can be built.
Our model directly builds on the Toner-Bacon model, which simulates maximally entangled states using one bit of communication and shared randomness toner . By selecting an instance at random, our model reproduces the same statistics as the single-world model in Ref. gisin , in which outcomes are generated through shared randomness and the single use of a nonlocal PR box.
In Sec. II, we discuss the general framework of a single-world ontological theory. In Sec. III, the framework is extended to the case of multiple coexisting realities. This extension includes many-worlds theory as limiting case. In Sec. IV, a simplified version of DH theory is discussed in the two-party Bell scenario. In Sec. V, we introduce our model with finite information flow. A discussion follows in Sec. VI. Finally, we draw the conclusions and perspectives. In particular, we remark that the introduced framework offers a generalized setting for quantum communication complexity, which has its own appeal beyond foundational motivations.
II Single-world ontological theories
Ontological theories of quantum mechanics have been motivated by the apparent contradiction between the well-defined macroscopic reality and the linear evolution of the wave function, leading to the measurement problem. Since the cat in Schrödinger’s thought experiment must be dead or alive after all, some auxiliary information about his actual state should be added to the information contained in the wave function, provided that the linearity of the Schrödinger equation is assumed universal. This completion is consistently realized, for example, in the dBB theory bohm , in which both the wave function and the position of the particles are part of the ontology. In this respect, the dBB theory could be considered a kind of many-worlds theory brown with extra-variables labeling which branch is experienced. Is any completion of quantum theory just a many-worlds theory augmented with extra-variables? That is, should the wave function be considered a physically real entity in any ontological theory? In the context of quantum foundations, such completions, typically referred to as $\psi$-ontic, are opposed to $\psi$-epistemic theories, where the wave function represents only an observer’s knowledge or information about a system. This distinction is extensively discussed in the seminal paper by N. Harrigan and R. W. Spekkens harrigan , who advocate for the $\psi$-epistemic view. Let us quote the closing lines of their abstract: “[…] Einstein was seeking not just any completion of quantum theory, but one wherein quantum states are solely representative of our knowledge”. In particular, the authors point out that one of Einstein’s arguments for incompleteness of quantum mechanics leads to support the epistemic view of quantum states under the premise of local causality. Unfortunately, Einstein’s argument alone is not supportive of the $\psi$-epistemic view, since local causality contradicts quantum mechanics under the condition of single actual outcomes, as shown by Bell. However, there are other reasons for promoting a $\psi$-epistemic view spekkens . Furthermore, $\psi$-epistemic models have an important role in quantum communication complexity, since they are essentially equivalent to classical simulations of quantum communication that employ finite classical communication montina_comm_epist0 ; montina_comm_epist . Since the $\psi$-epistemic view is the only possible alternative to an ontological theory which is not just many-worlds theory with extra-variables, it merits further consideration.
An ontological theory of quantum phenomena is characterized by an ontological space $\Omega$ of ontic states, denoted by $\lambda$, which are interpreted as physically real states of the system. Generally, ontological models are introduced in the context of a quantum system $S$, which undergoes a preparation and a subsequent measurement harrigan . In that context, the ontic state refers to the quantum system $S$, whereas measurements are described by indicator functions, which are conditional probabilities of getting an outcome given $\lambda$. As our focus is on a theory rather than a model, we include macroscopic systems interacting with $S$ in the description. Thus, the ontic state contains, in particular, full information about macroscopic facts, such as the position of the pointer of a measuring device. For example, in Schrödinger’s thought experiment, there exists a subset $\Omega_{A}\subset\Omega$ of the ontological space such that $\lambda\in\Omega_{A}$ if and only if the cat is alive.
Ontic states are typically assumed to evolve deterministically, or more generally, according to some causal rule, such as a Markov process. Since quantum theory is fundamentally probabilistic, the preparation of a quantum state $\psi$ does not uniquely determine the ontic state $\lambda$, but rather specifies a probability distribution $\rho(\lambda|\psi)$, possibly conditioned on additional contextual parameters, that is, we have the mapping
$\psi\rightarrow\rho(\lambda|\psi).$ (1)
Crucially, the converse is also true in $\psi$-epistemic ontological theories, that is, the ontic state $\lambda$ does not uniquely determine the wave function. This is equivalent to state that the wave function is not an element of the ontology. Within this framework, the branching of the wave function is not necessarily a physical process, since the wave function itself is not taken to represent a physical entity. Instead, its role is epistemic: its information is statistically encoded in the distribution of the underlying ontic state. Suggesting that a wave function branches into multiple realities would, in this view, be akin to claiming that a tossed coin splits into two worlds merely because each outcome is possible. Thus, a $\psi$-epistemic ontological theory is not a many-worlds theory with extra-variables. The coexistence of multiple realities will be an additional hypothesis justified by Bell nonlocality, rather than the branching of the wave function.
III Ontological theories with multiple coexisting realities
Curiously, the term $\psi$-epistemic has traditionally been used only in the context of single-world theories. In contrast, within the many-worlds community, the view that the wave function has ontological significance appears to be ubiquitous. For example, in Ref. kent , the ontological status of $\psi$ is identified as one of the two essential characteristics of the galaxy of many-worlds interpretations. Kent states: “ DeWitt (writes) ‘[…] The symbols of quantum mechanics represent reality just as much as do those of classical mechanics.’ There seems to be no dispute in the literature on this point: if a theory is not mathematically realist then it is not a many-worlds interpretation.” Indeed, the ontology of the wave function motivates the branching into many parallel worlds. As far as we know, the epistemic view has never been considered in a scenario involving multiple coexisting realities. Remarkably, the model presented in Ref. hall replaces the wave function with many interacting worlds, so that the quantum state is not treated as an element of reality in the theory. However, the model – which is nonlocal – reproduces quantum mechanics exactly only in the limit of infinitely many worlds. In this limit, the collective classical state of the infinite worlds contains complete information about the quantum state. Thus, the theory becomes $\psi$-ontic as the discrepancy with quantum predictions vanishes.
In this section, we extend the framework of single-world ontological theories by assuming that multiple realities may coexist. These theories are necessarily probabilistic as long as the number of realities is kept finite. This contrasts with the many-worlds theory, where probabilities must be justified or postulated a posteriori. The determinism of many-worlds theory can be recovered in the limit of infinite realities. As previously said, if the considered framework is $\psi$-epistemic, then the branching is not motivated by the branching of the wave function, rather by Bell’s theorem. Our aim is to explore whether quantum correlations can be understood in a locally causal framework, simply by relaxing the assumption of a single realized world without adopting the DH many-worlds theory as a whole.
This section is organized as follows. In Sec. III.1, we extend the single-world ontological theories by introducing multiple coexisting instances of some macroscopic physical quantity. In Sec. III.2, we clarify the terms instances and branches in our context by introducing an alternative representation using occupation numbers, which is more reminiscent of the formalism of many-worlds theory, in particular of Saunders’ branch counting. Furthermore, we show that many-worlds theory can be recovered (with some extra structure) in the limit of infinite instances. In Sec. III.3, the combination of two macroscopic quantities is discussed, which is preparatory for the definition of separability and locality of the theory, given in Sec. III.4.
III.1 Multiple instances of a macroscopic quantity
[…]
$\lambda=(s_{1},s_{2},\dots,s_{m},\eta),$ (2)
[…]
Since the preparation of the quantum state does not uniquely determine the ontic state $\lambda$, the $m$-tuple of outcomes $(s_{1},s_{2},\dots,s_{m})$ generally emerges from the measurement with probability $\rho(s_{1},s_{2},\dots,s_{m})$. Like in many-worlds theory, a single observer will experience one of the $m$ outcomes. In a single-world theory, one would have a distribution $\rho(s)$ over a single outcome $s$, consistent with Born’s rule. In our $m$-instances scenario, however, we must introduce an additional assumption to connect the joint distribution $\rho(s_{1},s_{2},\dots,s_{m})$ with the empirical data observed by a single observer:
Assumption (Equiprobability of instances). The observer is equally likely to be in either reality, without any weight attached to the instances.
This is motivated by the principle of indifference (also known as the principle of insufficient reason). The same assumption has been made in Refs. hall ; brassard . Accordingly, the probability of observing outcome $s$ is
$\rho(s)=\frac{1}{m}\sum_{k}\sum_{s_{1},\dots,s_{m}}\delta_{s_{k},s}\rho(s_{1},\dots,s_{m})$ (3)
This is equivalent to a simple unweighted counting rule: we count how many of the $m$ realities yield outcome $s$ (from 0 to $m$ realities), average over many runs, and divide by $m$. This evaluated $\rho(s)$ must agree with Born’s rule.
Two points are crucial here. First, the number of instances with given outcome $s$ fluctuates from run to run, such that the observed frequencies converge to those predicted by quantum theory as the number of runs goes to infinity. If the theory were deterministic, one would observe only frequencies equal to a rational number in $\{0,1/m,2/m,\dots,1\}$ (the latter with $s=s_{1}=s_{2}=s_{3}=\dots=s_{m}$). Second, in each individual run, not all outcomes need to be realized, in contrast to the many-worlds framework, where all the branches of the wave function are instantiated. In our example, at most $m$ distinct outcomes can occur in a single run (only one if $s_{1}=s_{2}=\dots=s_{m}$). If $m<n$, where $n$ is the number of branches of the wave function, we would have at most $m$ distinct outcomes per run, not the $n$ outcomes realized in the many-worlds theory. Even if $m>n$, the number of distinct realities may be smaller than $n$ in some runs of the experiment. In particular, if there exists a branch of the wave function with an extremely small probability $p$, this means that, based on our knowledge of the system, there is an extremely small probability (not greater than $mp$) that some instance realizes that branch. Thus, it is not guaranteed that such an instance exists in every run of the experiment. This is trivially true in a single-world ontological theory with only one existing reality. This departs sharply from the deterministic realization of all the branches in many-worlds theories. Provided that the ontic state $\lambda$ does not contain $\psi$, the framework also departs from dBB theory, where there are branches which are physically real, but empty of observers.
It is trivial to prove that this framework is consistent with quantum mechanics. In the de Broglie-Bohm theory with one instance ($m=1$), $\eta$ is identified with the wave function and $s_{1}$ represents the position of all the particles in the laboratory (including the particles of the pointer). A naive many-instances theory can be trivially built by introducing $m$ trajectories in the pilot-wave theory such that the position of the particles in each instance is distributed according to Born’s rule. Whereas such many-instances extensions of the de Broglie-Bohm theory shows that the framework does not conflict with quantum mechanics, they have no appealing motivation. The dBB theory consistently solves the measurement problem (which was the motivation of the theory), but raises new problems (nonlocality bell ) and critics (existence of “empty” branches brown ). Introducing additional trajectories does not weaken these new concerns. It is decisively more interesting to investigate $\psi$-epistemic theories of multiple instances, in which $\eta$ is not identified with the quantum state or, more generally, $\lambda$ does not completely define the quantum state.
At this point, it may be useful to clarify the terms instances, realities and branches used in our context. This will further highlight the similarities between our general framework and many-worlds theory – the latter being a limiting case of the former. Later, we will discuss how to combine different macroscopic quantities. This will allow us to introduce the concepts of separability and locality.
III.2 Branches versus instances
Branches, worlds, and realities are interchangeable terms in the context of Everett’s many-worlds theory. When the wave function goes into a superposition of macroscopically distinct states, we say that one world branches into separate worlds or macroscopically distinct realities. Thus, by definition, two branches, worlds or realities differ by some physical quantity. For example, if the physical quantity has $n$ different values and each value has a nonzero probability amplitude, we say that we have $n$ different branches or worlds, although they may have different weights.
In our context, there is a distinction between realities, or instances, and the branches identified by the values taken by physical quantities. Instances are meant as realities with their own private ontic variables $s_{k}$, but these variables may take identical values in different instances. Consequently, we say that the system branches into two distinguishable realities when some physical quantity in different instances evolves from one shared initial value to two distinct values. Let us further clarify this point by discussing an alternative representation of the ontological theory, which aligns more closely with the narrative of the many-worlds theory.
[…]
$\lambda=(n_{1},\dots,n_{M},\eta),$ (4)
[…]
The new representation using occupation numbers is more akin to the usual picture of a many-worlds theory. The label $i$ identifies a branch, whereas the occupation numbers, which are integers, replace the quantum probabilities weighting the branches of many-worlds theory. Provided that the number of branches exceeds the total occupation number, it is clear that many branches have zero occupation number (zero instances). In a single-world ontological theory, the total occupation number is just one, so that only one branch is realized.
Our weighting of branches with an occupation number also somehow resembles Saunders’ branch counting saunders . Saunders argues that the quantum states continuously undergo a huge number of branching at the microscopic level. These microscopically distinguishable states result into few indistinguishable macroscopic states. By partitioning macroscopic branches into suitable sets of microscopic branches with equal weight, Born’s rule is recovered through a simple count of microscopic branches in the limit of infinite partitioning. Saunders’ microstates and macrostates are akin to instances and branches in our terminology. As in our case, each value of a macroscopic observable is weighted by an integer. The mark difference is that Saunders’ weights must go to infinity to recover Born’s rule because of determinism. Conversely, the weights $n_{i}$, which are stochastic variables, are not proportional to the quantum probabilities in a single run – Born’s rule is recovered only on average. This is what happens in every single-world ontological theory, in which only one value has occupation number equal to $1$, whereas the others are equal to zero. The value taken by the single instance changes stochastically from run to run. Consider, for example, the de Broglie-Bohm theory, in which the particle positions are probabilistically distributed according to Born’s rule.
III.2.1 Many-worlds theory in the limit $m\rightarrow\infty$
[…]
$\bar{n}_{i}=|\psi_{i}|^{2}$ (5)
in each run. The quantity $\bar{n}_{i}$ is the fraction of instances in the branch $i$, which resembles Saunders’ branch counting. This quantity offers a convenient way to explain why branches with greater weights are more likely to be experienced – they are associated with a higher multitude of instances, so that the previously stated assumption of equiprobability of instances holds. In the limit $N_{t}\rightarrow\infty$, the framework with constraint (5) is de facto equivalent to many-worlds theory with the quantities $\bar{n}_{i}$ as extra structure. The constraint (5) may be justified by the diverse derivations of Born’s rule given in many-worlds theory. For example, the constraint can be derived by assuming that $\bar{n}_{i}$ is a function of $|\psi_{i}|$ and that an additivity property holds, along the same lines as Everett everett . Namely, the additivity property states that the probability weight of $\sum_{k}a_{k}|\phi_{k}\rangle$ is equal to the sum of the probability weights of $a_{1},a_{2},\dots$, where $|\phi_{1}\rangle,|\phi_{2}\rangle,\dots$ are orthonormal vectors. This property guarantees consistency of the constraint under unitary evolutions – namely, it ensures the normalization of the probabilities over time.
It can be interesting to investigate whether constraint (5) can be imposed only on average by retaining the stochasticity of the framework in the limit $N_{t}\rightarrow\infty$.
III.2.2 Conservation of the number of instances
The representation using occupation numbers might raise the question of whether the total number of instances can be created or annihilated, analogous to particles in quantum field theory. Given the analogy between occupation numbers and the quantum probabilities weighting the branches, the conservation of the total number of instances appears as a natural hypothesis, mirroring the conservation of the norm of the wave function. Indeed, the creation of new instances would correspond to the ex novo creation of macroscopic systems – such as observers and measuring devices. This would be peculiar, especially considering that each instance is regarded as a physically existing system. Such creation would imply an increase in energy. Indeed, branching in the many-worlds theory does not involve the creation of new systems, but rather a divergence into different states of the same system. Similarly, in our framework, branching corresponds to a fixed number of instances whose states diverge from a shared initial state.
III.3 Combination of two macroscopic quantities
Before introducing separability and locality in the theory, let us discuss the rule for combining two macroscopic quantities, say ${\cal M}_{A}$ and ${\cal M}_{B}$, generated by performing two different measurements on the quantum system $\cal Q$}. Each of these quantities are, for example, the pointer of some measuring device. There are different ways of generating ${\cal M}_{A}$ and ${\cal M}_{B}$. Let us analyze three cases:
  1. 1.
    One observer, say Alice, performs first one measurement generating the macroscopic quantity ${\cal M}_{A}$, then she performs a second measurement generating ${\cal M}_{B}$.
  2. 2.
    Alice performs a measurement, generating ${\cal M}_{A}$, then departs from the measured system. Subsequently, another party, Bob, performs a second measurement, generating ${\cal M}_{B}$. Alice and Bob have not interacted with each other in the interval between the measurements.
  3. 3.
    Given two entangled systems ${\cal Q}_{A}$ and ${\cal Q}_{B}$ that are spacelike-separated, Alice and Bob perform a local measurement on ${\cal Q}_{A}$ and ${\cal Q}_{B}$, respectively. Eventually, they meet to compare their results.
Let us assume that there are $m$ realities of Alice. In the first case, when Alice performs the first measurement, the pointer of one measuring device in the $k$-th instance moves to the position $s_{A,k}$, which is observed by Alice. After a second measurement is performed, the pointer of another device moves to position $s_{B,k}$, which is again observed. Since the number of realities of Alice is preserved, in the end, the overall quantity ${\cal M}={\cal M}_{A}\oplus{\cal M}_{B}$ has $m$ instances with value $(s_{A,k},s_{B,k})$ in the $k$-th instance. Thus, the value $s_{k}$ in Eq. (2) is identified with the pair $(s_{A,k},s_{B,k})$. It is worth to note that the combination of the quantities corresponds to a direct sum, rather than a tensor product.
The second case is more intricate. Let us assume that Bob has a number $m^{\prime}$ of instances. Since Alice and Bob never interact in the interval between the measurements, all we can say immediately after the last measurement is that Alice has $m$ instances of $s_{A}$ and Bob $m^{\prime}$ instances of $s_{B}$. There is no reason for pairing their respective outcomes as long as they stay far away. The ontic state is
$\lambda=(s_{A,1},s_{A,2},\dots,s_{A,m},s_{B,1},s_{B,2},\dots,s_{B,m^{\prime}},\eta),$ (6)
in which the ordering of $s_{A}$ and $s_{B}$ establishes no pairing rule of Alice and Bob’s instances. In this respect, Peres’ dictum ”unperformed measurements have no result” peres holds. Only when Alice and Bob meet, a pairing of their instances is defined. Since the number of instances of each party is preserved and one party eventually experiences one definite macroscopic state of the other party, the pairing must be one-to-one, so that $m^{\prime}=m$. For example, if Alice had fewer instances, some of them would need to pair with more than one of Bob’s. Such a mismatch would generally require the creation of Alice’s instances.
The specific way in which the pairing occurs depends on the ontic state at the moment the meeting takes place. The pairing is understood as a physical process in which one instance of Alice interacts with only one instance of Bob. Note that if the theory is deterministic, the ontic state contains information about the pairing even prior to the meeting of the parties. It is akin to saying that a neutron and a proton will bind if they are brought sufficiently close together. We seek a local mechanism for the pairing, which, however, may be mathematically defined in a nonlocal way at any given time.
The framework may be generalized to infinite instances. In this generalization, it is required that the cardinality of the set of instances is identical for each quantity (e.g. countable, continuum and so on).
Summarizing, we have the following properties
  • Each macroscopic quantity has a fixed number $m$ of instances (more generally, the set of instances has fixed cardinality).
  • If two quantities ${\cal Q}_{A}$ and ${\cal Q}_{B}$ are jointly known, then there is a one-to-one correspondence between the instances of ${\cal Q}_{A}$ and the instances of ${\cal Q}_{B}$. This correspondence defines a combination ${\cal Q}_{A}\oplus{\cal Q}_{B}$ with $m$ instances. The correspondence is determined by the ontic state $\lambda$.
The third case can be treated in the same way and is discussed in the following subsection.
III.4 Separability and locality
Until now, we have introduced a general framework of an ontological theory with multiple realities. As previously said, the generalization to multiple instances is suggested by Bell’s theorem, which is proved under the assumption of single actual outcomes. Thus, we are particularly interested to formulate a separable and local framework. Here, we will consider the Bell scenario of two spacelike-separated laboratories (Case 3 of Sec. III.3). Each laboratory includes a system (${\cal Q}_{A}$ and ${\cal Q}_{B}$), a measuring device, and an observer – Alice and Bob. The device has a pointer that is set in a ready state and indicates an outcome $s$ after a measurement is performed on the system. The two systems ${\cal Q}_{A}$ and ${\cal Q}_{B}$ are generally in some entangled state. Eventually, Alice and Bob compare the results at a meeting point. Before the measurements, the instances are macroscopically indistinguishable, so that there is only one macroscopically distinct realization of Alice and Bob. As previously said, to avoid the creation of new instances over time, we assume that each spatially distinct object has the same number $m$ of instances.
Separability – Like in a separable one-world ontological theory, the two laboratories have their own ontic states, say
$\left\{\begin{array}[]{l}\lambda_{A}=(s_{A,1},s_{A,2},\dots,s_{A,m},\eta_{A}),\vskip 2.84526pt\\ \lambda_{B}=(s_{B,1},s_{B,2},\dots,s_{B,m},\eta_{B}).\end{array}\right$. (7)
It is important to emphasize again that the $m\times m$ possible combinations of local instances do not correspond to $m\times m$ instances of the overall system composed of the two laboratories. Each party has exactly $m$ instances at all times – including when they meet to compare their results. At the meeting point, one instance of Alice will interact with exactly one instance of Bob, and vice versa. Thus, given the one-to-one pairing rule, there will be $m$ pairs of Alice and Bob after they interact.
Locality – Initially, at time $t=0$, the ontic states are $\lambda_{A}^{\prime}=(s_{A,1}^{\prime},s_{A,2}^{\prime},\dots,s_{A,m}^{\prime},\eta_{A}^{\prime})$ and $\lambda_{B}^{\prime}=(s_{B,1}^{\prime},s_{B,2}^{\prime},\dots,s_{B,m}^{\prime},\eta_{B}^{\prime})$, $s_{A,k}^{\prime}$ and $s_{B,k}^{\prime}$ being in the ready state 0. Later, Alice and Bob perform each a measurement randomly selected among a set of possibilities. Let us denote by parameters $a$ and $b$ their respective selection. As usual in Bell’s scenario, we assume that $a$ and $b$ are not correlated with $\lambda_{A}^{\prime}$ and $\lambda_{B}^{\prime}$. These parameters are randomly generated outside the laboratories and passed to Alice and Bob. The ontic states evolve to $\lambda_{A}^{\prime\prime}=(s_{A,1}^{\prime\prime},s_{A,2}^{\prime\prime},\dots,s_{A,m}^{\prime\prime},\eta_{A}^{\prime\prime})$ and $\lambda_{B}^{\prime\prime}=(s_{B,1}^{\prime\prime},s_{B,2}^{\prime\prime},\dots,s_{B,m}^{\prime\prime},\eta_{B}^{\prime\prime})$, where $s_{A,k}^{\prime\prime}$ and $s_{B,k}^{\prime\prime}$ take the value of the outcome in the $k$-th instance of Alice and Bob, respectively.
At this point, we assume that the theory is locally causal bell76 . On Alice side we impose that the probability of getting $\lambda_{A}^{\prime\prime}$ given $a$, $b$, $\lambda_{A}^{\prime}$, $\lambda_{B}^{\prime}$ and $\lambda_{B}^{\prime\prime}$ does not depend on $b$, $\lambda_{B}^{\prime}$ and $\lambda_{B}^{\prime\prime}$. Similarly, we impose a specular condition on Bob side. That is,
$\begin{array}[]{l}\rho(\lambda_{A}^{\prime\prime}|\lambda_{B}^{\prime\prime},\lambda_{A}^{\prime},\lambda_{B}^{\prime},a,b)=\rho(\lambda_{A}^{\prime\prime}|\lambda_{A}^{\prime},a)\vskip 2.84526pt\\ \rho(\lambda_{B}^{\prime\prime}|\lambda_{A}^{\prime\prime},\lambda_{A}^{\prime},\lambda_{B}^{\prime},a,b)=\rho(\lambda_{B}^{\prime\prime}|\lambda_{B}^{\prime},b).\end{array}$ (8)
These conditions lead to the factorization
$\rho(\lambda_{A}^{\prime\prime},\lambda_{B}^{\prime\prime}|\lambda_{A}^{\prime},\lambda_{B}^{\prime},a,b)=\rho(\lambda_{A}^{\prime\prime}|\lambda_{A}^{\prime},a)\rho(\lambda_{B}^{\prime\prime}|\lambda_{B}^{\prime},b),$ (9)
from which Bell’s inequalities are derived in the single-world case bell76 . Generally, $\lambda_{A}^{\prime}$ and $\lambda_{B}^{\prime}$ are correlated, as the systems are entangled. Under the hypothesis of determinism, the assumption of local causality can be replaced by the weaker assumption of locality. See Ref. wiseman for a comprehensive discussion on the conceptual differences between the two assumptions in the context of Bell’s proofs.
Comparing the outcomes – After the measurements, Alice and Bob reach the meeting point and compare their results. The $i$-th instance of Alice is paired with the $k_{i}$-th instance of Bob, where $k_{i}=k_{i}(\lambda_{A}^{\prime\prime},\lambda_{B}^{\prime\prime})$ is some bijective function in $i$ depending on the ontic states $\lambda_{A}^{\prime\prime}$ and $\lambda_{B}^{\prime\prime}$. Depending on the model, the pairing may need only a partial information on $a$ and $b$, which is encoded in $\lambda_{A}^{\prime\prime}$ and $\lambda_{B}^{\prime\prime}$. Conversely, the DH theory needs the full information on the setting parameters (as shown later in Sec. IV). In Sec. V, we will see that the correlation of two maximally entangled qubits can be reproduced by a local model in which the two parties need to carry just one bit of information about their performed measurement to the meeting point. Furthermore, the model requires just two instances per party. In general, the number of realities and the amount of information can depend on the dimension of the Hilbert space of each subsystem, the number of subsystems and the computational network within which the subsystems interact.
In the next section, we will introduce a simplified version of the DH theory applied to the two-party Bell scenario, then we will introduce the $\psi$-epistemic model.
IV Nonlocal correlations in many-worlds theory
In Ref. deutsch , Deutsch and Hayden proposed a local, many-worlds interpretation of quantum correlations by introducing a set of operators – referred to as descriptors – that evolve in the Heisenberg picture. This framework significantly expands the ontology of the theory by incorporating these new entities. Here, we present a highly simplified caricature of their argument. While this oversimplification may make the argument seem somewhat trivial, it nevertheless captures its essence. We consider Bell’s scenario with two local von Neumann measurements on two entangled qubits. Rather than using the descriptors, we describe each qubit by the operational parameters defining the measurements which are performed on them.
Suppose that two qubits are prepared in the singlet state
$|\Psi\rangle=\frac{1}{\sqrt{2}}\left(|-1,1\rangle-|1,-1\rangle\right).$
The qubits are sent to two spatially separated parties, Alice and Bob. Each party independently chooses to perform a projective measurement, represented by the Bloch vectors $\vec{a}$ and $\vec{b}$, respectively. Alice gets $s_{A}=\pm 1$ and Bob $s_{B}=\pm 1$. The outcomes of each measurement occur with equal probability, that is,
$P(s_{A}|\vec{a})=P(s_{B}|\vec{b})=1/2.$ (10)
In the DH framework, this means that the qubits locally split into two branches with equal weights. On each side, this process occurs without the need to know what is happening on the other side, as a consequence of the no signaling property of quantum theory. At this point of the experiment, no weight is attached on the pair of outcomes $(s_{A},s_{B})$. After the measurements, the parties decide to compare their results. To do that, they have to meet at the same place or send some communication through a classical channel. After multiple running of the experiment over different entangled states, they find out that the outcomes, $s_{A}\in\{-1,1\}$ and $s_{B}\in\{-1,1\}$, are distributed according to the probability distribution
$P(s_{A},s_{B}|\vec{a},\vec{b})=\frac{1}{4}\left(1-s_{A}s_{B}\vec{a}\cdot\vec{b}\right),$ (11)
which violate the Bell inequalities. While this violation leads to a break of local realism in a single-world scenario, the DH framework evades this conclusion as follows. When Alice meets with Bob, the branch in which she observed $s_{A}$ further divides into two branches labeled by the value $s_{B}$ observed by Bob. The probability of being paired with Bob observing $s_{B}$ given her observation $s_{A}$ is
$P(s_{B}|s_{A},\vec{a},\vec{b})\equiv P(s_{A},s_{B}|\vec{a},\vec{b})/P(s_{A}|\vec{a}).$ (12)
Thus, at the meeting point, four branches emerge such that each one is weighted with an amplitude whose modulus square is exactly $P(s_{A},s_{B}|\vec{a},\vec{b})$. By construction, these weights are not in conflict with the previous weight assignments. Rephrasing in Saunders’ terms, the (infinite) multitudes of microstates of each party are paired in such a way that the fraction of realities experiencing the outcomes $s_{A}$ and $s_{B}$ is equal to $P(s_{A},s_{B}|\vec{a},\vec{b})$. All these processes occur locally, which is guaranteed by the no signaling conditions
$\begin{array}[]{l}\sum_{s_{B}}P(s_{A},s_{B}|\vec{a},\vec{b})=P(s_{A}|\vec{a})\\ \sum_{s_{A}}P(s_{A},s_{B}|\vec{a},\vec{b})=P(s_{B}|\vec{b}).\end{array}$ (13)
It is worth recalling again Peres’ dictum that “unperformed measurements have no result”. The meeting is like a new measurement in which the pair of outcomes $(s_{A},s_{B})$ is made definite. As long as the comparison of the results does not occur, the parties can only claim their results (in their own reality), but the pairing of the two local realities remains indefinite. The idea of postponing the build-up of the correlations is illustrated in Ref. brassard for nonlocal PR boxes and further discussed in the general quantum case in Ref. charles . The model in Ref. brassard simplifies the pairing rule at the meeting point as the emerging branches have no weight. In particular, the rule is identical to the rule discussed in Sec. III.4: each instance of one party is paired with one instance of the other party.
Thus, the existence of many local realities allows us to delay the buildup of the correlations until the parties meet or communicate each other. It is fundamental to remark that, in the DH theory, each party has to carry the information about the performed measurement at the meeting point in order to get the right pairing. Namely, the probability $P(s_{A},s_{B}|\vec{a},\vec{b})$ must be known, which cannot be evaluated without knowing the Bloch vectors $\vec{a}$ and $\vec{b}$. Thus, each party has to communicate two real parameters, that is, infinite information. Employing the descriptors of Ref. deutsch , the number of real parameters would be $4^{2}-4=12$ per party charles2 . The total number of independent parameters is $15$, since the descriptors of Alice are not completely independent of the descriptors of Bob. This infinite amount of information is a consequence of the $\psi$-ontic nature of the theory.
V A model with finite information flow
Employing the framework of Sec. III.4, let us show that two unweighted instances and a finite information flow are enough for simulating the outcomes of all the projective measurements $\vec{a}$ and $\vec{b}$ on a maximally entangled state. The proof is a simple adaptation of the Toner-Bacon model toner to the ontological model of Sec. III.4 with $m=2$ instances per party. The Toner-Bacon model provides a classical simulation of the outcomes by using one bit of communication between the parties and some shared randomness. Let us first review this single-world model.
The classical protocol is as follows. Alice and Bob share two random and independent unit vectors, $\vec{x}_{0}$ and $\vec{x}_{1}$, uniformly distributed over a sphere. First, Alice chooses to simulate the measurement $\vec{a}$ and generates an outcome $s_{A}$ such that
$s_{A}=\text{sign}(\vec{a}\cdot\vec{x}_{0}).$ (14)
Then, she generates a number
$n_{A}=\text{sign}(\vec{a}\cdot\vec{x}_{0})\text{sign}(\vec{a}\cdot\vec{x}_{1}),$ (15)
which is sent to Bob. Finally, Bob chooses to simulate the measurements $\vec{b}$ and generates the outcome
$s_{B}=-\text{sign}\left[\vec{b}\cdot(\vec{x}_{0}+n_{A}\vec{x}_{1})\right].$ (16)
Averaging on $\vec{x}_{0}$ and $\vec{x}_{1}$, it turns out that the outcomes are generated according to Eq. (11). Let us reproduce the proof of Ref. toner by using the Kochen-Specker model kochen as shortcut in the last step. It is sufficient to show that
$\langle s_{A}s_{B}\rangle=-\vec{a}\cdot\vec{b}.$ (17)
We have
$\begin{array}[]{l}\langle s_{A}s_{B}\rangle=\frac{1}{(4\pi)^{2}}\sum_{n_{A}=\pm 1}\int d^{2}x_{0}d^{2}x_{1}\text{sign}(\vec{a}\cdot\vec{x}_{0})\vskip 2.84526pt\\ \text{sign}\left[-\vec{b}\cdot(\vec{x}_{0}+n_{A}\vec{x}_{1})\right]\frac{1+n_{A}\text{sign}(\vec{a}\cdot\vec{x}_{0})(\vec{a}\cdot\vec{x}_{1})}{2}.\end{array}$ (18)
First, we note that the two terms of the sum over $n_{A}$ are identical (replace $\vec{x}_{1}$ with $\vec{x}_{1}$ in the second term). Using the distributive property and a suitable swapping $\vec{x}_{0}\leftrightarrow\vec{x}_{1}$ , we have
$\langle s_{A}s_{B}\rangle=\frac{1}{8\pi^{2}}\int d^{2}x_{0}d^{2}x_{1}\text{sign}(\vec{a}\cdot\vec{x}_{0})\text{sign}\left[-\vec{b}\cdot(\vec{x}_{0}+\vec{x}_{1})\right].$ (19)
Integrating over $\vec{x}_{1}$ in spherical coordinates with $\vec{b}$ as the pole, we have
$\langle s_{A}s_{B}\rangle=\frac{1}{2\pi}\int d^{2}x_{0}(-\vec{b}\cdot\vec{x}_{0})\text{sign}(\vec{a}\cdot\vec{x}_{0}),$ (20)
which can be recast in the form
$\langle s_{A}s_{B}\rangle=\frac{1}{\pi}\int d^{2}x_{0}(-\vec{b}\cdot\vec{x}_{0})\theta(\vec{b}\cdot\vec{x}_{0})\theta(\vec{a}\cdot\vec{x}_{0})-(\vec{a}\rightarrow-\vec{a})$ (21)
where $\theta(x)$ is the Heaviside function. The first term appears in the Kochen-Specker model kochen and is $(1+\vec{a}\cdot\vec{b})/2$. Thus, Eq. (17) is proved.
Now, let us reinterpret this model within the framework of Sec. III.4 with $m=2$ instances per party. The resulting model is schematically represented in Fig. 1. The variables $\eta_{A}$ and $\eta_{B}$ in Sec. III.4 contain the integers $n_{A}$ and $n_{B}$, respectively, and a copy of the random variables $\vec{x}_{0}$ and $\vec{x}_{1}$. Initially, the two instances are indistinguishable. When the measurements are performed, each party branches into two distinct alternatives, labeled $A_{\pm 1}$ and $B_{\pm 1}$ for Alice and Bob, respectively. The two parties are not allowed to communicate until they reach the meeting point, where each alternative of Alice is paired to one alternative of Bob according to some rule that we will determine. First, we note that the Toner-Bacon model generates the same correlation if $s_{A}$ and $s_{B}$ are both chosen with opposite sign. Let us assume that alternatives $A_{1}$ and $A_{-1}$ get the outcome $s_{A}$ in Eq. (14) and its opposite, respectively. They both set $n_{A}$ according to Eq. (15). All these processes involve only local information that is available to Alice. Thus, the ontic state $\lambda_{A}$ of Eq. (7) evolves locally according to Eq. (8). A complication arises in deciding which outcomes are associated with Bob’s alternatives. Associating $s_{B}$ in Eq. (16) to alternative $B_{1}$ would require to know the value $n_{A}$, an information that is not available to Bob. If we set $n_{A}$ equal to $1$ and, finally, we pair alternative $A_{w}$ with $B_{w}$ at the meeting point, we end up making a mistake in the case $n_{A}$ were 1-1 and $\text{sign}\left[\vec{b}\cdot\vec{x}_{+}\right]\neq\text{sign}\left[\vec{b}\cdot\vec{x}_{-}\right]$, where $\vec{x}_{\pm}\equiv\vec{x}_{0}\pm\vec{x}_{1}$. Crucially, the mistake can be corrected during the pairing by knowing just one classical bit of information per measurement, namely $n_{A}$ for Alice’s measurement and
$n_{B}\equiv\text{sign}(\vec{b}\cdot\vec{x}_{+})\text{sign}(\vec{b}\cdot\vec{x}_{-})$ (22)
for Bob’s measurement. Thus, let us assume that alternative $B_{1}$ and $B_{-1}$ generate the outcomes
$s_{B}=-\text{sign}\left[\vec{b}\cdot(\vec{x}_{0}+\vec{x}_{1})\right]$ (23)
and $s_{B}$, respectively. Furthermore, both Bob’s alternatives generate $n_{B}$ according to Eq. (22), which will be used to define the pairing. Also the ontic state $\lambda_{B}$ of Eq. (7) evolves locally according to Eq. (8). Thus, the overall model is local. Finally, Bob and Alice reach the meeting point to compare their results. Bob and Alice carry information about their outcome and the numbers $n_{A}$ and $n_{B}$, respectively. Alternatives $A_{\pm 1}$ are paired with $B_{\pm 1}$, unless $n_{A}=n_{B}=-1$, in which case the pairing is swapped. This leads to the following pairing rule
$\begin{array}[]{c}(n_{A},n_{B})\neq(-1,-1)\Rightarrow A_{\pm 1}\leftrightarrow B_{\pm 1}\\ (n_{A},n_{B})=(-1,-1)\Rightarrow A_{\pm 1}\leftrightarrow B_{\mp 1}.\end{array}$ (24)
Let us stress that there are only two instances after the pairing. If $(n_{A},n_{B})\neq(-1,-1)$, then Alice and Bob agree that the pair of outcomes is $(s_{A},s_{B})$ in one instance and $(-s_{A},-s_{B})$ in the other instance. Conversely, if $(n_{A},n_{B})=(-1,-1)$, the pairing is flipped and the instances are $(s_{A},-s_{B})$ and $(-s_{A},s_{B})$. It is worth to note that only $2$ of the $4$ possible combinations of outcomes are realized in each run. If there were $4$ unweighted instances deterministically realizing all the combinations, the probability of observing one pair of outcomes would be $1/4$ in each run. To match quantum theory using a finite number of unweighted instances, the final pair of outcomes must be stochastic – that is, they must vary from run to run. From the perspective of one realization of Alice and Bob, the probability of getting a pair of outcomes $(s^{\prime},s^{\prime\prime})\in\{-1,1\}^{2}$ is obtained by a simple count of unweighted instances. First, we compute the fraction of instances getting $(s^{\prime},s^{\prime\prime})$. Then, we average over $\vec{x}_{0}$ and $\vec{x}_{1}$. There is no amplitude weight of the branches.
[…]
Figure 1: Local model simulating correlations of local projective measurements on two maximally entangled qubits. Alice and Bob receive each two random vectors $\vec{x}_{0}$ and $\vec{x}_{1}$ from a common source. Given projective measurements $\vec{a}$ and $\vec{b}$, they locally generate $\{n_{A},\vec{s}_{A}\equiv(s_{A},-s_{A})\}$ and $\{n_{B},\vec{s}_{B}\equiv(s_{B},-s_{B})\}$, respectively, by using only locally available information. The numbers $s_{A}$, $n_{A}$, $s_{B}$ and $n_{B}$ are generated according to Eqs. (14152322), respectively. Finally, they meet at a common point and each instance of one party is paired with one instance of the other party according to the rule in the figure. This results in two instances of the overall outcomes.
Summarizing, the model is as follows.
The $\psi$-epistemic local model.
Simulation of local von Neumann measurements with two instances and one communicated bit per party in Bell’s scenario of two maximally entangled qubits (Fig. 1).
  • Alice and Bob receive the random vectors $\vec{x}_{0}$ and $\vec{x}_{1}$.
  • In instance $A_{\pm 1}$, Alice generates outcome $\pm s_{A}$ according to Eq. (14), given the measurement setting $\vec{a}$. She also sets $n_{A}$ by using Eq. (15).
  • In instance $B_{\pm 1}$, Bob generates outcome $\pm s_{B}$ according to Eq. (23), given the measurement setting $\vec{b}$. He sets $n_{B}$ by using Eq. (22).
  • Finally, they reach the meeting point. Instance $A_{\pm 1}$ is paired with instance $B_{\pm 1}$ if $(n_{A},n_{B})\neq(-1,-1)$, otherwise the pairing is swapped.
At each run of the simulation, two instances of the overall outcomes are generated. The quantum probabilities are obtained by selecting one of the two instances at random with equal probability (assumption of equiprobability of instances in Sec. III.1).
It is interesting to note that the pairing depends only on $n_{A}$ and $n_{B}$. There is no need to carry the initial random vectors $\vec{x}_{0}$ and $\vec{x}_{1}$, which are used only for generating $s_{A}$, $s_{B}$, $n_{A}$ and $n_{B}$.
This model is mathematically equivalent to the model in Ref. gisin , but they differ in the interpretation. The latter is a single-world model in which the parties share a nonlocal PR box as a resource for generating the outcome. In our model, we have a “local” PR box which is used for pairing locally the alternatives at the meeting point. The statistics of the single-world model of Ref. gisin is recovered by randomly taking one of the two paired alternatives in the two-instances model. It is surprising that a simple and seemingly trivial conceptual step separates our model from that in Ref. gisin , yet it has gone unnoticed until now.
It is interesting to note that a local single-world model could be constructed in a slightly different scenario. Instead of generating the correct outcomes at the moment measurements are performed, the task shifts to generating them at the meeting point. Each party only needs to communicate two bits. However, including the whole process in this local description would lead to the paradoxical conclusion that a party’s state – and even their memory of past observations – can suddenly change. In a many-instances framework, this macroscopic “rewind” and memory reconstruction is avoided by allowing two distinct realities to evolve independently.
VI Discussion
The nonlocality arising in the framework of single-world ontological theories can lead the proponents of Deutsch-Hayden many-worlds theory to claim the failure of a probabilistic framework in which $\psi$ represents a mere state of knowledge. The main purpose of this paper is to show with a proof of principle that the coexistence of multiple realities is sufficient for a local account of the correlations of two maximally entangled qubits. There is no known proof that both determinism and the ontology of the quantum states are necessary conditions for a local theory. Furthermore, the drop of these two assumptions can lead to a more “economical” theory, in which the information flow is made as small as possible and the quantum probabilities are obtained by a simple count of unweighted instances.
Let us discuss in detail the two key differences between the model in Sec. V and the Deutsch-Hayden many-worlds theory – namely, the way the probabilities come out and the role played by the quantum state.
Probabilities – In the scenario simulated by the model, if Alice and Bob perform the measurements $\vec{a}$ and $\vec{b}$, then they observe outcome $s_{A}$ and $s_{B}$ with probability $(1-s_{A}s_{B}\vec{a}\cdot\vec{b})/4$. The many-worlds theory says that there are $4$ branches, one for each outcome, weighted by an amplitude whose modulus square is the observed probability. In Saunders’ terms, there is a multitude of branches of microstates such that the ratio between the number of branches with outcomes $s_{A}$ and $s_{B}$ and the number of overall branches is equal to the observed probability. Since this probability is a real number, this multitude is actually infinite.
In our model, the story is quite different. There is a finite number of unweighted instances (namely, two instances). The observed probability of the final pair of outcomes $(s_{A},s_{B})$ is obtained by taking the fraction of instances with outcome $(s_{A},s_{B})$ (the fraction is 0 or $1/2$ in the specific model) and averaging over different runs. Like in the general framework of Sec. III, the different instances have equal probability of being experienced by an observer. Note that some outcomes are not realized in a single run of the experiment (in MW theory, all the outcomes are realized in each run). Indeed, two instances cannot realize all the $4$ possible combinations of outcomes. The remarkable point of the model is not the reduction from $4$ to $2$ realized branches, but the equiprobabilities of the two instances. This is possible because of the stochasticity of the model. As previously said, the derivation of the branch weights of many-worlds theory from a branch counting à la Saunders saunders would require a division of each of the $4$ branches into an infinite multitude of equally probable instances (see also Sec. III.2.1, in which it is shown that determinism is recovered in the limit of infinite instances in our framework).
Recalling the general discussion of Sec. III, let us transliterate this feature of the model to an experiment generating the superposition
$\psi_{+}|+\rangle+\psi_{-}|$-$\rangle$ (25)
where $|\pm\rangle$ are two macroscopically distinct states, and $|\psi_{-}|^{2}=10^{-40}|\psi_{+}|^{2}$. In many-worlds theory, both the alternatives are realized, however we expect to experience $|+\rangle$ with near-certainty. In our framework, we have a finite number $m$ of instances such that the ratio between the number of instances associated with state $|$-$\rangle$ and the number of all the instances is equal to our expectation of experiencing $|$-$\rangle$ in a single run of the experiment. If the total number of instances is small ($m\ll 10^{40}$), this ratio is zero in most of the runs. That is, the branch $|$-$\rangle$ is almost never realized. The ratio turns out to be equal to $|\psi_{-}|^{2}$ only on average over a large number of runs (like in any single-world ontological theory). This point of view is in the middle between a single-world ontological theory, such as de Broglie-Bohm theory, and many-worlds theory. In the former, there is some ontic variable saying that the system is always realized in one of the two superposed states. The latter says that both the states are always realized in parallel worlds. As a single-world theory is nonlocal, as proved by Bell, we may need more than one world for a local description, but between 11 (single-world ontological theories) and infinity (many-worlds theory plus Saunders’ branch counting) there is any integer greater than 11!
Role of the quantum state – The second difference is more subtle. In a $\psi$-epistemic theory, the ontic state does not contain the full information about the quantum state. Rather, this information is encoded in the statistics of the ontic state. Let us consider the quantum state emerging from Alice’s measurement and how its information is encoded in Deutsch-Hayden theory and our model. In the former, the ongoing state, which has to be carried to the meeting point, is the vector $\vec{a}$ and the outcome associated with each branch. Their knowledge allows us to infer the outgoing quantum state after the measurement in each branch, meaning that the quantum state is part of the ontology. In our model, this inference is not possible in a single run of the experiment. After the measurement, Alice generates a number $n_{A}$ and associates an outcome to each branch. This information and the knowledge of the shared random variables $\vec{x}_{0}$ and $\vec{x}_{1}$ enable us only to infer that the quantum state $\vec{a}$ in branch $A_{1}$ after the measurement is any unit vector such that Eqs. (14,15) are fulfilled. Rather, the full information about quantum state is statistically encoded in the ontic variables. This is necessary for making the information flow from the parties to the meeting point finite montina_comm_epist0 ; montina_comm_epist . It is worth to recall again that the whole ontic state after the measurements is not actually required at the meeting point, since $\vec{x}_{0}$ and $\vec{x}_{1}$ are not involved in the pairing rule.
VII Conclusions
In conclusion, we have introduced a general probabilistic framework that encompasses both single-world and many-worlds theories as limiting cases. This hybrid approach, which combines the randomness of single-world ontological theories with the multiplicity of coexisting realities found in many-worlds theory, offers a way to address the challenges inherent in each individual previous framework.
  • On one hand, the coexistence of diverse realities allows us to evade Bell’s theorem, which applies to single-world ontological theories. More broadly, it may also circumvent the contextuality requirement. For instance, a manifestation of preparation contextuality is the proof that ontological theories inherently break time symmetry leifer1 . However, with appropriate adaptations, our model provides a fully time-symmetric description of a scenario in which a qubit, in a maximally mixed state, is measured at two different times monti_erase . The time asymmetry proved in Ref. leifer1 is transferred to the measurement devices and the subsequent comparison of results, which are inherently time-asymmetric processes. Similarly, the branching circumvents the recently proved theorem that measurements erase information monti_erase , which is another manifestation of contextuality. An open question is whether a many-instances framework can also elude the Pusey-Barrett-Rudolph (PBR) theorem pbr , which implies that quantum mechanics is in conflict with $\psi$-epistemic theories under a single-world premise and an assumption of preparation independence (a kind of locality condition). Note that this bypass of the theorem has never been considered so far, since Deutsch-Hayden theory is local but also $\psi$-ontic. It is also interesting to note that the hypothesis of preparation independence may be questioned in the framework of nonlocal single-world ontological theories, but it becomes mandatory in our framework, which aims to provide a locally causal structure. One possible way to circumvent the theorem is to assume that multiple preparations may coexist, which is more drastic than assuming that multiple outcomes coexist.
  • On the other hand, the randomness of the framework makes it possible to recover Born’s rule from a simple count over a finite set of instances. Furthermore, it can lead to a reduction of the information flow as much as possible by stripping the quantum state of its “ontic” rank. Finally, if the quantum state represents our knowledge about a system rather than an element of reality, then it is possible that not all alternatives in a superposition are necessarily realized (as discussed in Sec. III.1 and illustrated by our model). If, according to prior knowledge, there existed a branch with a very faint amplitude in which Scipio was defeated in the Third Punic War, this would not necessarily imply that there is a world where Carthage rules over Rome. That is, the multiple instances of the framework discussed in Sec. III.1 may not realize all possible branches.
Here, we have presented a proof of principle of this hybridization by showing that local measurements on two maximally entangled qubits can be locally simulated with only two unweighted instances, finite communication and shared randomness. It is plausible that a generalization to many qubits and any quantum state will require a blow-up of the shared variables. This increase of local resources occurs also in the Deutsch-Hayden many-worlds theory deutsch , in which the descriptors are represented by matrices acting on the whole Hilbert space of all involved particles. The main difference, however, is that the shared random variables in our approach would not depend on the state of the system – they would be generated before the “game” starts and would provide a stochastic background over which the system evolves. A similar idea has been suggested in Ref. montina_non_markov within a single-world scenario. Our purpose has been to introduce a general framework with a simple proof of principle, rather than a comprehensive theory. Whether this hybrid approach can be extended to a general local theory distinct from DH theory remains an open and, in our opinion, fascinating question. A negative answer to this question would suggest that every locally causal theory of quantum mechanics is essentially the DH theory.
A generalization of the presented model to higher dimensional systems can be obtained from the approximate model in Ref. montina_approx which extends the Toner-Bacon protocol to arbitrary dimensions. Although this generalized model is not exact, its error is very small (about $1\%$) for few qubits per party. Its adaptation to our many-instances framework would provide a very accurate model for a sufficiently low dimension. This adaptation can suggest improvements for reducing the error or finding exact protocols.
From a broader perspective, this framework provides a generalized setting for quantum communication complexity. Rather than just quantifying the communication cost of classically simulating quantum entanglement, one can consider a scenario involving a third party – a referee – who receives multiple alternative outcomes and determines how to pair them. In this setting, the question concerns not only the minimal communication cost from the two parties to the referee, but also the minimal number of parallel unweighted instances required for a local simulation of quantum correlations. The standard communication complexity problem of quantifying the required communication cost for simulating entanglement is a subclass of this generalized setting with the slight variation that the generation of one of the outcomes is postponed and delegated to the referee. The subclass corresponds to the situation in which there is one instance and one of the two parties – say Bob – communicates to the referee the full information about his choice. Let us underline that, in this case, Bob is not supposed to generate the outcome, but delegates the task to the referee. A further generalization could allow for multiple parallel choices of measurements performed by Alice and Bob. Focusing only on one side of the problem, one can consider $\psi$-ontic models with infinite communication and find the minimal number of instances for an exact local simulation. It can be interesting, for example, to investigate some kind of ‘pilot’ theory with a finite number of instances in which each instance is locally piloted by Deutsch-Hayden descriptors (rather than the wave-function like in pilot-wave theories). The task would be to find the minimal number of instances and the pairing rules that lead to an exact simulation of quantum correlations for more general entangled states.
In the context of this generalized setting in quantum communication complexity, there is an interesting aspect that is worth mentioning and offers perspectives for future research. On one hand, it is known that the communication cost of simulating quantum correlations goes to zero as the entangled state approaches a product state renner . Thus, if we adopt the communication cost as a measure of ‘nonlocality’, weakly entangled states are less ‘nonlocal’ than strongly entangled states (here, ‘nonlocality’ refers to single-world scenarios). On the other hand, maximally entangled states can be simulated using a single ‘nonlocal’ PR-box gisin , whereas weakly entangled states require more than one PR-box brunner . Thus, adopting the number of PR-boxes as a measure of nonlocality, we obtain the opposite conclusion that weakly entangled states are more ‘nonlocal’ than maximally entangled states. In our generalized setting of communication complexity, the task is not just to evaluate the minimal communication cost, but also the minimal number of parallel realities (instances) for a local account of quantum correlations. It is entirely possible that these two quantities behave differently as functions of the ‘degree’ of entanglement. In particular, the number of required instances is likely related to the second measure of ‘nonlocality’ (with PR-boxes). Thus, a local model with multiple instances could potentially ‘explain’ why weakly entangled states are more robust against the detection loophole in an experimental Bell test loophole . Indeed, if more instances are required for a correct simulation, a single-world protocol is less likely to succeed by discarding some simulation runs instead of correctly pairing the instances. Thus, our framework is not just relevant in quantum foundations, but suggests new technical problems which have their own appeal for classical simulations of quantum entanglement and quantum communication. Our framework may provide new insights into some of the counter-intuitive behaviors of entanglement.
Another line of future research can be inspired by the approach of many interacting worlds in Ref. hall , which has been built on ground of the dBB theory. As in our framework, this approach employs a finite number of worlds but is equivalent to quantum mechanics only in the limit of infinite worlds. Thus, it is experimentally testable. The approach is nonlocal, so that one can wonder whether a modification making it “local” is possible. For example, in the search of such a modified approach, one could start from the descriptors, rather than the wave function.
Acknowledgments
We wish to thank Charles Bédard for useful comments and suggestions. We also thank Lev Vaidman for pointing out that some statements in the previous version are not accepted by the entire Everettian community. Finally, we are grateful to Sophie Berthelette for her careful reading of the manuscript. The authors acknowledge support from the Swiss National Science Foundation (SNSF), project No. 200020- 214808.